# Geo-Sign: Hyperbolic Contrastive Regularisation for Geometrically Aware Sign Language Translation

### v1

@@ 4.3 Qualitative Analysis: Visualizing Embedding Spaces
To intuitively understand the effect of hyperbolic regularisation, we visualise the learned pose embeddings. Figure 2 shows UMAP [43] projections of these embeddings into the 2D Poincaré disk (by log-mapping hyperbolic embeddings to the tangent space at the origin, then applying UMAP). We compare embeddings from our Geo-Sign (Hyperbolic Token) model against those from the Geo-Sign (Euclidean Token) model, which uses the same contrastive token-level alignment but without hyperbolic projection (curvature $c=0$).
@@
The Euclidean embeddings (Figure 2, Left) appear relatively clustered and undifferentiated. In contrast, the hyperbolic embeddings (Figure 2, Right) exhibit a more structured distribution. Notably, embeddings corresponding to hand articulations (often carrying fine-grained lexical information) tend to occupy regions further from the origin, towards the periphery of the Poincaré disk. This is consistent with hyperbolic geometry’s property of expanding space near the boundary, providing more capacity to distinguish subtle variations. Conversely, features representing larger body movements or overall posture (often conveying prosodic or grammatical information) tend to be located more centrally. This visualised structure suggests that the hyperbolic model indeed learns to place features in a manner that reflects the hierarchical nature of sign kinematics, with fine details pushed to high-curvature regions and global features remaining near the low-curvature origin. This geometric organization likely contributes to the improved discriminability and translation performance observed.
@@
Figure 2: UMAP projection of pose part summary embeddings ($\bar{\mathbf{f}}_{p}$ onto the 2D Poincaré disk). (Left) Embeddings from the Euclidean Token regularisation model ($c=0.001$). (Right) Embeddings from the Geo-Sign (Hyperbolic Token) model. The hyperbolic embeddings show a more structured distribution, with hand features (representing finer details) often pushed towards the periphery, while body/face features (representing coarser semantics) are more central, indicative of a learned kinematic hierarchy.
5 Conclusion and Future Work
@@
This paper introduced Geo-Sign, a novel framework that enhances Sign Language Translation by leveraging hyperbolic geometry to model the inherent hierarchical structure of sign language kinematics. By projecting skeletal features from ST-GCNs into the Poincaré ball and employing a geometric contrastive loss, Geo-Sign regularises a pre-trained mT5 model, guiding it to learn more discriminative and geometrically aware representations. We explored two alignment strategies: a global pooled method and a fine-grained token-based attention method operating directly in hyperbolic space. Our experimental results on the CSL-Daily benchmark demonstrate the significant benefits of this approach.
5.1 Limitations
The quality of skeletal representations remains dependent on upstream pose estimation. While offering representational benefits, hyperbolic operations can add computational overhead compared to purely Euclidean ones, though this is generally offset by avoiding raw video processing. The optimal choice of hyperbolic model parameters (e.g., curvature strategy) warrants further study. Generalizability to a wider range of sign languages also needs investigation.
5.2 Future Work
Promising directions include exploring other hyperbolic models (e.g., Lorentz), developing more sophisticated dynamic curvature adaptation, integrating Geo-Sign’s hyperbolic skeletal features into multi-modal frameworks, and applying these geometric principles to other sign language processing tasks like recognition or generation. Further research into the interpretability of learned hyperbolic embeddings could also yield deeper insights into how sign language structure is captured.
@@ Appendix A Introduction
  • The specifics of pose feature extraction and the Spatio-Temporal Graph Convolutional Network (ST-GCN) architecture employed (Section C.1).
  • Detailed explanations and implementations of our proposed hyperbolic alignment strategies, including the Pooled Method and the Token Method (Section C.2).
  • Further mathematical derivations and discussions pertinent to hyperbolic operations, such as Fréchet mean computation and contrastive loss gradients (Appendix D).
  • Elaboration on the learnable parameters within our model, particularly the manifold curvature $c$ and the loss blending factor $\alpha$ (Appendix E).
  • A discussion of computational considerations, experimental setup, and qualitative results (Appendix F).
  • Key code snippets for essential components of Geo-Sign are provided in Appendix G to aid in understanding and replication.
@@ C.1.2 ST-GCN Architecture
    Graph Definition: The skeletal structure for each part is defined as a graph, where joints are nodes and natural bone connections are edges. The Graph class, detailed in LABEL:lst:gcn_utils_graph (Appendix G), handles the construction of these graphs and their corresponding adjacency matrices.
@@
    ST-GCN Blocks: A sequence of ST-GCN blocks processes these features. Each block (see STGCN_block in LABEL:lst:stgcn_block_chain, Appendix G) consists of:
@@
    Residual Connections: To allow richer feature interaction, residual connections are introduced from the body stream’s ST-GCN output to the hand and face streams before their final temporal fusion layers. This allows global body posture context to inform the interpretation of fine-grained hand and face movements. Details are in LABEL:lst:models_residual_gcn (Appendix G). This design choice treats body features as fixed contextual input for the parts during each forward pass, isolating the body feature extractor from direct updates via part-specific losses.
@@ C.2.1 Pooled Method (Global Semantic Alignment)
1. Part-Specific Hyperbolic Embeddings: The temporally mean-pooled Euclidean feature vectors $\bar{\mathbf{f}}_{p}$ for each anatomical part $p$ (body, hands, face) are projected into the Poincaré ball $\mathbb{B}_{c}^{d_{\text{hyp}}}$. This projection, yielding hyperbolic embeddings $\mathbf{h}_{p}$, is achieved using the HyperbolicProjection layer (LABEL:lst:hyperbolic_projection in Appendix G), as defined in Eq. (4) of the main paper:
@@
Here, $\lambda_{w}$ is a temperature parameter for the softmax (e.g., fixed to 1.0 in our experiments) controlling the sharpness of the weight distribution. The computation is performed iteratively as detailed in Algorithm 1 of the main paper and LABEL:lst:frechet_mean (Appendix G).
@@
The implementation details are shown in LABEL:lst:pooled_text_emb (Appendix G).
@@ C.2.2 Token Method (Fine-Grained Part-Text Alignment)
3. Hyperbolic Attention Mechanism: For each hyperbolic pose part embedding $\mathbf{h}_{p}$ (acting as a query), a contextual text embedding $\mathbf{c}_{p}$ is generated. This is achieved using a hyperbolic attention mechanism (see LABEL:lst:token_attention in Appendix G) that operates as follows:
  • Key Transformation: The hyperbolic text token embeddings $\{\mathbf{h}_{\text{token},j}\}$ serve as keys. The embeddings are first transformed using learnable Möbius transformations to enhance their representational capacity: k_j = (M_key ⊗_c h_token,j) ⊕_c b_key, where $\mathbf{M}_{\text{key}}$ is a learnable Möbius matrix and $\mathbf{b}_{\text{key}}$ is a learnable Möbius bias vector.
  • Attention Scores: Attention scores are computed based on the negative geodesic distance between each pose query $\mathbf{h}_{p}$ and each transformed text key $\mathbf{k}_{j}$: score_pj = -d_B_c(h_p, k_j).
  • Attention Weights: These scores are normalised using a softmax function (after applying padding masks) to obtain attention weights $\alpha_{pj}$: α_pj = softmax(scorepjτattn), where $\tau_{\text{attn}}$ is a learnable temperature parameter for the attention mechanism, distinct from the temperature in the contrastive loss.
@@ C.2.3 Intuition Behind the Token Method
    • Standard attention weights $\alpha_{pj}$ then quantify the contribution of each text token $j$ to the meaning conveyed by pose part $p$.
@@ D.1 Fréchet Mean in the Poincaré Ball
Why distance-based weights?  Each pose part (body, face, left hand, right hand) yields a hyperbolic embedding $h_{p}$. We set $w_{p}\propto\exp\!\bigl{(}d_{\mathbb{B}_{c}}(0,h_{p})/\lambda_{w}\bigr{)}$ so parts farther from the origin—in regions of higher curvature and greater discriminative power—receive more influence. Without this weighting the mean would drift toward the centre, diluting information contributed by the hands and face.
@@ D.2 Gradient of the Hyperbolic Distance
[…]
@@ E.1 Discussion on Learnable Curvature
The curvature of the Poincaré ball, $\kappa=-c$ (where $c>0$), is a crucial hyperparameter that dictates the “shape” of the hyperbolic space. Instead of fixing $c$ heuristically, we make it a learnable parameter of our model (see LABEL:lst:manifold_init in Appendix G).
@@
The main paper’s ablation studies (Table 2a) show that initializing $c$ in the range of $1.0-2.0$ (e.g., optimal BLEU-4 at $c=1.5$) yields strong performance. Figure 3 illustrates how $c$ adapts during training from different initializations.
@@
Figure 3: Evolution of the learnable manifold curvature $c$ during training for different initializations. (2(a)) When initialised at $c=1.50$, the curvature magnitude slightly decreases, suggesting an optimal value around $1.42$ for this setup. (2(b)) When initialised at a low $c=0.10$, the curvature increases, indicating the model benefits from more “hyperbolic space” initially. It stabilizes around $c=0.20$, potentially influenced by the dynamic $\alpha$ schedule that reduces regularization emphasis over time.
@@ E.2 Discussion on Loss Blending Factor $\alpha$
The total training loss $\mathcal{L}_{\text{total}}$ is a weighted combination of the primary cross-entropy translation loss $\mathcal{L}_{\text{CE}}$ and our hyperbolic contrastive regularization term $\mathcal{L}_{\text{hyp\_reg}}$: L_total = αL_CE + (1-α) ⋅L_hyp_reg. The blending factor $\alpha$ is not fixed but is dynamically adjusted during training. This dynamic scheduling allows the model to potentially benefit from different loss emphases at different training stages. The calculation of $\alpha$ at each training step (see LABEL:lst:alpha_calc in Appendix G) is:
@@ F.1.1 Profiler Summary and Comparative Analysis
Table 5 summarizes key metrics from the profiler. Parameter counts are consistent with the main paper’s Table 1, while MACs (Multiply-Accumulate operations) and Latency are derived from DeepSpeed profiler outputs for a batch size of 8. Section F.1.1 provides a comparison of model parameters against other gloss-free methods.
@@
Baseline Uni-Sign (Pose) $587.75$ $116.59$ $415.73$
@@
Geo-Sign (Hyperbolic Token) $589.10$ $1.35$ $116.60$approx9.96 $2550.00$ $513.4$
@@
In both Geo-Sign variants, the parameter overhead from hyperbolic components is less than $0.25\%$ of the total model size. As shown in Section F.1.1, our Geo-Sign models achieve competitive or superior performance to recent RGB-based methods while maintaining a significantly smaller total parameter count. This highlights the efficiency of enhancing skeletal representations with geometric priors, challenging the trend that relies solely on scaling up visual encoders and language model decoders for performance gains in SLT.
@@
The MACs from the learnable linear transformations within the hyperbolic projection layers constitute a very small fraction ($<0.01\%$) of the total model MACs. The bulk of MACs originates from the mT5 model (profiled at $\approx 66.29$ GMACs) and the ST-GCN modules (profiled at $\approx 49.93$ GMACs). We should note, however, that standard profilers (like DeepSpeed’s MAC counter) primarily quantify MACs from common operations like convolutions and linear layers. The computational cost of specialised geometric functions within geoopt (e.g., manifold.dist, expmap0, logmap0, Möbius arithmetic) is not explicitly broken out as distinct hyperbolic operation MACs. These functions often involve sequences of elementary operations that are not all MAC-based (e.g., square roots, divisions, trigonometric functions like artanh or tanh). Thus, their computational load may be underestimated by MAC counters and is often better reflected in measured latency.
@@
  • Geo-Sign (Hyperbolic Pooled): $\approx 1630\text{ ms}$ (1.63 s), an increase of $\approx 1214\text{ ms}$ or $\approx 292\%$ over the baseline (approx. 3.9×3.9$\times$3.9 × slowdown).
  • Geo-Sign (Hyperbolic Token): $\approx 2550\text{ ms}$ (2.55 s), an increase of $\approx 2134\text{ ms}$ or $\approx 513\%$ over the baseline (approx. 6.1×6.1$\times$6.1 × slowdown).
@@ F.2 Further Technical Implementation Details
  • Hyperparameter Tuning Strategy: Key hyperparameters specific to the hyperbolic components, such as the initial curvature $c$, the initial loss blending factor $\alpha_{\text{init}}$ (referred to as args.alpha in code/main paper), and the hyperbolic embedding dimension $d_{\text{hyp}}$, were tuned using a grid search strategy on the CSL-Daily development set. Full hyperparameters are outlined in Table 7.
@@
    • Tangent Vector Clipping: Before applying an exponential map $\exp_{\mathbf{x}}^{c}(\mathbf{v})$ from a point $\mathbf{x}$ with a tangent vector $\mathbf{v}$, especially $\exp_{\mathbf{0}}^{c}(\mathbf{v})$, it’s crucial to ensure the resulting point remains strictly within the Poincaré ball and that the norm of $\mathbf{v}$ doesn’t cause numerical issues in $\tanh(\cdot)$. We apply a clipping strategy as mentioned in Section 3.4 of the main paper: v_clippedvmax(1, c∥v2+ ϵclip), for a small $\epsilon_{\text{clip}}>0$ (e.g., $10^{-5}$). This ensures that the argument to $\tanh$ in $\exp_{\mathbf{0}}^{c}$ does not become excessively large and that mapped points do not reach or exceed the boundary of the Poincaré ball. The project=True flag in geoopt’s expmap functions also helps enforce this by projecting points back onto the ball if they numerically fall outside.
@@
In Table 7 we provide the full hyper-parameters for the best performing model. The full code will be released following the review process.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
F.3 Qualitative Results
Additional Figures: Figure 4 (similar to aspects shown in Figure 2 of the main paper, concerning learned embedding distributions) illustrates the dynamic utilization of the hyperbolic manifold by showing the average geodesic distance of different pose part embeddings from the origin during training. Notably, features corresponding to hand articulations, which often carry fine-grained lexical information, tend to migrate towards the periphery of the Poincaré disk. This suggests that the model leverages the increased representational capacity in high-curvature regions to distinguish subtle hand-based signs.
Furthermore, Figure 5 (again, related to Figure 2 of the main paper, specifically the UMAP projections) provides a PCA-reduced visualization of the learned hyperbolic pose part embeddings projected onto the 2D Poincaré disk for 1000 poses. This plot reveals a structured distribution where body features cluster near the origin (a more Euclidean-like region suitable for broader semantics), while hand and face features are more dispersed, with hand features populating regions further towards the boundary. This geometric organization, reflecting a learned kinematic hierarchy, likely contributes to the improved discriminability and, consequently, the enhanced translation quality demonstrated in the following examples. These visualizations support the hypothesis that the geometric biases induced by hyperbolic space aid in forming more effective representations for sign language translation.

### v2

@@ 4.3 Qualitative Analysis: Visualizing Embedding Spaces
To intuitively understand the effect of hyperbolic regularisation, we visualise the learned pose embeddings. Figure˜2 shows UMAP [47] projections of these embeddings into the 2D Poincaré disk (by log-mapping hyperbolic embeddings to the tangent space at the origin, then applying UMAP). We compare embeddings from our Geo-Sign (Hyperbolic Token) model against those from the Geo-Sign (Euclidean Token) model, which uses the same contrastive token-level alignment but without hyperbolic projection (curvature $c=0$).
@@
The Euclidean embeddings (Figure˜2, Left) appear relatively clustered and undifferentiated. In contrast, the hyperbolic embeddings (Figure˜2, Right) exhibit a more structured distribution. Notably, embeddings corresponding to hand articulations (often carrying fine-grained lexical information) tend to occupy regions further from the origin, towards the periphery of the Poincaré disk. This is consistent with hyperbolic geometry’s property of expanding space near the boundary, providing more capacity to distinguish subtle variations. Conversely, features representing larger body movements or overall posture (often conveying prosodic or grammatical information) tend to be located more centrally. This visualised structure suggests that the hyperbolic model indeed learns to place features in a manner that reflects the hierarchical nature of sign kinematics, with fine details pushed to high-curvature regions and global features remaining near the low-curvature origin.
@@
Figure 2: UMAP projection of pose part summary embeddings ($\bar{\mathbf{f}}_{p}$ onto the 2D Poincaré disk). (Left) Embeddings from the Euclidean Token regularisation model ($c=0.001$). (Right) Embeddings from the Geo-Sign (Hyperbolic Token) model. The hyperbolic embeddings show a more structured distribution, with hand features (representing finer details) often pushed towards the periphery indicative of a learned kinematic hierarchy.
5 Conclusion
@@
This paper introduced Geo-Sign, a novel framework that enhances Sign Language Translation by leveraging hyperbolic geometry to model the inherent hierarchical structure of sign language kinematics. By projecting skeletal features from ST-GCNs into the Poincaré ball and employing a geometric contrastive loss, Geo-Sign regularises a pre-trained mT5 model, guiding it to learn more discriminative and geometrically aware representations. We explored two alignment strategies: a global pooled method and a fine-grained token-based attention method operating directly in hyperbolic space. Our experimental results on both Chinese Sign Language and American Sign Language demonstrate the significant benefits of this approach.
@@ Appendix A Introduction
  • The specifics of pose feature extraction and the Spatio-Temporal Graph Convolutional Network (ST-GCN) architecture employed (Section˜C.1).
  • Detailed explanations and implementations of our proposed hyperbolic alignment strategies, including the Pooled Method and the Token Method (Section˜C.2).
  • Further mathematical derivations and discussions pertinent to hyperbolic operations, such as Fréchet mean computation and contrastive loss gradients (Appendix˜D).
  • Elaboration on the learnable parameters within our model, particularly the manifold curvature $c$ and the loss blending factor $\alpha$ (Appendix˜E).
  • A discussion of computational considerations, experimental setup, and qualitative results (Appendix˜F).
  • Key code snippets for essential components of Geo-Sign are provided in Appendix˜G to aid in understanding and replication.
@@ C.1.2 ST-GCN Architecture
    Graph Definition: The skeletal structure for each part is defined as a graph, where joints are nodes and natural bone connections are edges. The Graph class, detailed in LABEL:lst:gcn_utils_graph (Appendix˜G), handles the construction of these graphs and their corresponding adjacency matrices.
@@
    ST-GCN Blocks: A sequence of ST-GCN blocks processes these features. Each block (see STGCN_block in LABEL:lst:stgcn_block_chain, Appendix˜G) consists of:
@@
    Residual Connections: To allow richer feature interaction, residual connections are introduced from the body stream’s ST-GCN output to the hand and face streams before their final temporal fusion layers. This allows global body posture context to inform the interpretation of fine-grained hand and face movements. Details are in LABEL:lst:models_residual_gcn (Appendix˜G). This design choice treats body features as fixed contextual input for the parts during each forward pass, isolating the body feature extractor from direct updates via part-specific losses.
@@ C.2.1 Pooled Method (Global Semantic Alignment)
1. Part-Specific Hyperbolic Embeddings: The temporally mean-pooled Euclidean feature vectors $\bar{\mathbf{f}}_{p}$ for each anatomical part $p$ (body, hands, face) are projected into the Poincaré ball $\mathbb{B}_{c}^{d_{\text{hyp}}}$. This projection, yielding hyperbolic embeddings $\mathbf{h}_{p}$, is achieved using the HyperbolicProjection layer (LABEL:lst:hyperbolic_projection in Appendix˜G), as defined in Eq. (4) of the main paper:
@@
Here, $\lambda_{w}$ is a temperature parameter for the softmax (e.g., fixed to 1.0 in our experiments) controlling the sharpness of the weight distribution. The computation is performed iteratively as detailed in Algorithm 1 of the main paper and LABEL:lst:frechet_mean (Appendix˜G).
@@
The implementation details are shown in LABEL:lst:pooled_text_emb (Appendix˜G).
@@ C.2.2 Token Method (Fine-Grained Part-Text Alignment)
3. Hyperbolic Attention Mechanism: For each hyperbolic pose part embedding $\mathbf{h}_{p}$ (acting as a query), a contextual text embedding $\mathbf{c}_{p}$ is generated. This is achieved using a hyperbolic attention mechanism (see LABEL:lst:token_attention in Appendix˜G) that operates as follows:
  • Key Transformation: The hyperbolic text token embeddings $\{\mathbf{h}_{\text{token},j}\}$ serve as keys. The embeddings are first transformed using learnable Möbius transformations to enhance their representational capacity:
    $\mathbf{k}_{j}=(\mathbf{M}_{\text{key}}\otimes_{c}\mathbf{h}_{\text{token},j})\oplus_{c}\mathbf{b}_{\text{key}},$
    where $\mathbf{M}_{\text{key}}$ is a learnable Möbius matrix and $\mathbf{b}_{\text{key}}$ is a learnable Möbius bias vector.
  • Attention Scores: Attention scores are computed based on the negative geodesic distance between each pose query $\mathbf{h}_{p}$ and each transformed text key $\mathbf{k}_{j}$:
    $\text{score}_{pj}=-d_{\mathbb{B}_{c}}(\mathbf{h}_{p},\mathbf{k}_{j}).$
  • Attention Weights: These scores are normalised using a softmax function (after applying padding masks) to obtain attention weights $\alpha_{pj}$:
    $\alpha_{pj}=\text{softmax}\left(\frac{\text{score}_{pj}}{\tau_{\text{attn}}}\right),$
    where $\tau_{\text{attn}}$ is a learnable temperature parameter for the attention mechanism, distinct from the temperature in the contrastive loss.
@@ C.2.3 Intuition Behind the Token Method
    • The attention weights $\alpha_{pj}$ then quantify the contribution of each text token $j$ to the meaning conveyed by pose part $p$.
@@ D.1 Fréchet Mean in the Poincaré Ball
Why distance-based weights?  Each pose part (body, face, left hand, right hand) yields a hyperbolic embedding $h_{p}$. We set $w_{p}\propto\exp\!\bigl(d_{\mathbb{B}_{c}}(0,h_{p})/\lambda_{w}\bigr)$ so parts farther from the origin, in regions of higher curvature and greater discriminative power, receive more influence. Without this weighting the mean would drift toward the centre, diluting information contributed by the hands and face.
@@ D.2 Gradient of the Hyperbolic Distance
$\nabla_{u}\,d_{\mathbb{B}_{c}}(u,v)=-\frac{2}{\lambda_{u}^{c}\,\lambda_{v}^{c}}\,\frac{w}{\|w\|_{2}}\,\frac{1}{1-c\|w\|_{2}^{\,2}}$ (16)
@@ E.1 Discussion on Learnable Curvature
The curvature of the Poincaré ball, $\kappa=-c$ (where $c>0$), is a crucial hyperparameter that dictates the “shape” of the hyperbolic space. Instead of fixing $c$ heuristically, we make it a learnable parameter of our model (see LABEL:lst:manifold_init in Appendix˜G).
@@
The main paper’s ablation studies (Table 2a) show that initializing $c$ in the range of 1.02.01.0-2.0 (e.g., optimal BLEU-4 at $c=1.5$) yields strong performance. Figure˜3 illustrates how $c$ adapts during training from different initializations.
@@
Figure 3: Evolution of the learnable manifold curvature $c$ during training for different initializations. (3(a)) When initialised at $c=1.50$, the curvature magnitude slightly decreases, suggesting an optimal value around $1.42$ for this setup. (3(b)) When initialised at a low $c=0.10$, the curvature increases, indicating the model benefits from more “hyperbolic space” initially. It stabilizes around $c=0.20$, potentially influenced by the dynamic $\alpha$ schedule that reduces regularization emphasis over time.
@@ E.2 Discussion on Loss Blending Factor $\alpha$
The total training loss $\mathcal{L}_{\text{total}}$ is a weighted combination of the primary cross-entropy translation loss $\mathcal{L}_{\text{CE}}$ and our hyperbolic contrastive regularization term $\mathcal{L}_{\text{hyp\_reg}}$:
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{CE}}+(1-\alpha)\cdot\mathcal{L}_{\text{hyp\_reg}}.$
The blending factor $\alpha$ is not fixed but is dynamically adjusted during training. This dynamic scheduling allows the model to potentially benefit from different loss emphases at different training stages. The calculation of $\alpha$ at each training step (see LABEL:lst:alpha_calc in Appendix˜G) is:
@@ F.1.1 Profiler Summary and Comparative Analysis
Table˜5 summarizes key metrics from the profiler. Parameter counts are consistent with the main paper’s Table 1, while MACs (Multiply-Accumulate operations) and Latency are derived from DeepSpeed profiler outputs for a batch size of 8. Section˜F.1.1 provides a comparison of model parameters against other gloss-free methods.
@@
Baseline Uni-Sign (Pose) $587.75$ - $116.59$ - $415.73$ -
@@
Geo-Sign (Hyperbolic Token) $589.10$ $1.35$ $116.60$ $\approx$9.96 $2550.00$ $513.4$
@@
In both Geo-Sign variants, the parameter overhead from hyperbolic components is less than $0.25\%$ of the total model size. As shown in Section˜F.1.1, our Geo-Sign models achieve competitive or superior performance to recent RGB-based methods while maintaining a significantly smaller total parameter count. This highlights the efficiency of enhancing skeletal representations with geometric priors, challenging the trend that relies solely on scaling up visual encoders and language model decoders for performance gains in SLT.
@@
The MACs from the learnable linear transformations within the hyperbolic projection layers constitute a very small fraction ($0.01\%$) of the total model MACs. The bulk of MACs originates from the mT5 model (profiled at $\approx 66.29$ GMACs) and the ST-GCN modules (profiled at $\approx 49.93$ GMACs). We should note, however, that standard profilers (like DeepSpeed’s MAC counter) primarily quantify MACs from common operations like convolutions and linear layers. The computational cost of specialised geometric functions within geoopt (e.g., manifold.dist, expmap0, logmap0, Möbius arithmetic) is not explicitly broken out as distinct hyperbolic operation MACs. These functions often involve sequences of elementary operations that are not all MAC-based (e.g., square roots, divisions, trigonometric functions like artanh or tanh). Thus, their computational load may be underestimated by MAC counters and is often better reflected in measured latency.
@@
  • Geo-Sign (Hyperbolic Pooled): $\approx 1630\text{ ms}$ (1.63 s), an increase of $\approx 1214\text{ ms}$ or $\approx 292\%$ over the baseline (approx. $3.9\times$ slowdown).
  • Geo-Sign (Hyperbolic Token): $\approx 2550\text{ ms}$ (2.55 s), an increase of $\approx 2134\text{ ms}$ or $\approx 513\%$ over the baseline (approx. $6.1\times$ slowdown).
@@ F.2 Further Technical Implementation Details
  • Hyperparameter Tuning Strategy: Key hyperparameters specific to the hyperbolic components, such as the initial curvature $c$, the initial loss blending factor $\alpha_{\text{init}}$ (referred to as args.alpha in code/main paper), and the hyperbolic embedding dimension $d_{\text{hyp}}$, were tuned using a grid search strategy on the CSL-Daily development set. Full hyperparameters are outlined in Table˜7.
@@
    • Tangent Vector Clipping: Before applying an exponential map $\exp_{\mathbf{x}}^{c}(\mathbf{v})$ from a point $\mathbf{x}$ with a tangent vector $\mathbf{v}$, especially $\exp_{\mathbf{0}}^{c}(\mathbf{v})$, it’s crucial to ensure the resulting point remains strictly within the Poincaré ball and that the norm of $\mathbf{v}$ doesn’t cause numerical issues in $\tanh(\cdot)$. We apply a clipping strategy as mentioned in Section 3.4 of the main paper:
      $\mathbf{v}_{\text{clipped}}\leftarrow\frac{\mathbf{v}}{\max(1,\sqrt{c}\|\mathbf{v}\|_{2}+\epsilon_{\text{clip}})},$
      for a small $\epsilon_{\text{clip}}>0$ (e.g., $10^{-5}$). This ensures that the argument to $\tanh$ in $\exp_{\mathbf{0}}^{c}$ does not become excessively large and that mapped points do not reach or exceed the boundary of the Poincaré ball. The project=True flag in geoopt’s expmap functions also helps enforce this by projecting points back onto the ball if they numerically fall outside.
@@
In Table˜7 we provide the full hyper-parameters for the best performing model. The full code will be released following the review process.
@@
Gradient Accumulation Steps 8 Effective batch size becomes $8\times\text{accum\_steps}\times\text{num\_gpus}$
@@
Learning Rate (lr) $3\text{\times}{10}^{-5}$ For Euclidean parameters (AdamW)
@@
AdamW $\epsilon$ (opt-eps) $1\text{\times}{10}^{-8}$ Term for numerical stability
@@
Minimum LR (min-lr) $1\text{\times}{10}^{-6}$ Lower bound for LR in scheduler
@@
Learning Rate (hyp_lr) $1\text{\times}{10}^{-3}$ For hyperbolic parameters (RAdam)
@@
F.3 Limitations and Future Work
While offering representational benefits, hyperbolic operations can add computational overhead compared to purely Euclidean ones though this is generally offset by avoiding raw video processing. The optimal choice of hyperbolic model parameters (e.g., curvature strategy) warrants further study. Generalizability to a wider range of sign languages also needs investigation. Promising directions include exploring other hyperbolic models (e.g., Lorentz), developing more sophisticated dynamic curvature adaptation, integrating Geo-Sign’s hyperbolic skeletal features into multi-modal frameworks, and applying these geometric principles to other sign language processing tasks like recognition or generation. Further research into the interpretability of learned hyperbolic embeddings could also yield deeper insights into how sign language structure is captured.
F.4 Qualitative Results
Additional Figures: Figure˜4 (similar to aspects shown in Figure 2 of the main paper, concerning learned embedding distributions) illustrates the dynamic utilization of the hyperbolic manifold by showing the average geodesic distance of different pose part embeddings from the origin during training. Notably, features corresponding to hand articulations, which often carry fine-grained lexical information, tend to migrate towards the periphery of the Poincaré disk. This suggests that the model leverages the increased representational capacity in high-curvature regions to distinguish subtle hand-based signs.
Furthermore, Figure˜5 (again, related to Figure 2 of the main paper, specifically the UMAP projections) provides a PCA-reduced visualization of the learned hyperbolic pose part embeddings projected onto the 2D Poincaré disk for 1000 poses. This plot reveals a structured distribution where body features cluster near the origin (a more Euclidean-like region suitable for broader semantics), while hand and face features are more dispersed, with hand features populating regions further towards the boundary. This geometric organization, reflecting a learned kinematic hierarchy, likely contributes to the improved discriminability and, consequently, the enhanced translation quality demonstrated in the following examples. These visualizations support the hypothesis that the geometric biases induced by hyperbolic space aid in forming more effective representations for sign language translation.